OpenStudy (unofficialllyy):

Could someone help me with 5 Geometry questions? I will fan and medal :)

OpenStudy (unofficialllyy):

1. Given the triangle below, which of the following is a correct statement? \[\cot <B=\frac{ 6 }{ 5 }\] \[\csc <C=\frac{ 3 }{ 4 }\] \[\cot <C=\frac{ 8 }{ 5 }\] \[\csc <B=\frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

Do you still remember your inverse trigonometric ratios?

OpenStudy (unofficialllyy):

Yes. csc= hypotenuse/opposite sec=hypotenuse/adjacent cot=adjacent/opposite

OpenStudy (anonymous):

okay cool :) So from there, it is just trial and error as you see which of A,B,C or D is correct ^^

OpenStudy (unofficialllyy):

I'm kind of confused though, the csc options are 3/4 or 4/3 but where are they getting those numbers from?

OpenStudy (anonymous):

There are gotten, by interpreting the diagram :) Here I'll explain how to get the correct one !

OpenStudy (unofficialllyy):

I think it's the last option

OpenStudy (anonymous):

Yeah it is the last one !

OpenStudy (unofficialllyy):

I just realized that 8 is the hypotenuse and half of 8 is 4, and the opposite is 6 and half of 6 is 3 :)

OpenStudy (anonymous):

yeah, simplification is the way to go ^^

OpenStudy (unofficialllyy):

I have 4 more so let me post the next one. 2. Use △DEF, shown below, to answer the question that follows: What is the value of X rounded to the nearest hundredth? Type the numeric answer only.

OpenStudy (anonymous):

Ah no problem !

OpenStudy (anonymous):

So do you know how to start the next one? Do you remember your trig formulas for both Right angle and non-right angle triangles ?

OpenStudy (unofficialllyy):

No not off the top of my head

OpenStudy (anonymous):

Ah okay, I'll help out ! Sine Rule for non-right angle triangles is SinA = SinB a b

OpenStudy (anonymous):

Alternatively, when you forget next time, you can quickly search via the internet for the formulas : )

OpenStudy (anonymous):

In this particular question, we are working with the formula SinA = SinB a b

OpenStudy (unofficialllyy):

I think I got it, one sec.

OpenStudy (anonymous):

Alright cool ~

OpenStudy (unofficialllyy):

41.51?

OpenStudy (anonymous):

Yes ! that is correct

OpenStudy (unofficialllyy):

Yay!! 3. In △PQR, find the measure of ∡P. A. 30.4 degrees B. 35.9 degrees C. 54.1 degrees D. 59.6 degrees

OpenStudy (anonymous):

Alright, so now we are working with a right angle triangle, so the formulas we resort to are SOHCAHTOA

OpenStudy (anonymous):

Which of the standard trig ratios do you think it is ?

OpenStudy (unofficialllyy):

Do I use the tangent function ?

OpenStudy (anonymous):

Right again mate ^^

OpenStudy (unofficialllyy):

Okay I got 30.4

OpenStudy (anonymous):

Correct !

OpenStudy (unofficialllyy):

4. Find the value of the csc 40° using your calculator. This one is pretty easy lol I got 1.556

OpenStudy (anonymous):

yeap, way ahead of me haha, i just checked and your correct :)

OpenStudy (unofficialllyy):

5. In △JKL, solve for x. A. 74.89 B. 30.29 C. 38.16 D. 66.73

OpenStudy (anonymous):

By the way, do you sometimes forget your inverse trig ratios? Because I got a neat trick that you can't forget them :)

OpenStudy (unofficialllyy):

Yeah I do sometimes!

OpenStudy (anonymous):

The trick is to remember every 'third' letter of the inverse function For example Se'c' = Cos Co's'ec = Sin Co't' = Tan

OpenStudy (unofficialllyy):

Oh wow I didn't even realize that!

OpenStudy (anonymous):

Anyways, back on track :) I've got a solution for our next question ^^

OpenStudy (anonymous):

Yeah haha, its a hidden trick that I picked up a while back ~

OpenStudy (unofficialllyy):

Do I use the cosine function?

OpenStudy (anonymous):

yeah that's correct !

OpenStudy (unofficialllyy):

I got 38.16

OpenStudy (anonymous):

Yes that correct !

OpenStudy (unofficialllyy):

Thank you so much, you were so helpful!

OpenStudy (anonymous):

Glad to be of such help, anytime !